# Large Language Models for Software Engineering: A Systematic Literature Review

### v5

@@ 8.2.3. Enhancing LLM’s Performance in Existing SE Tasks
Tackling domain-specific challenges. Many SE domains, including safety-critical systems and specific industries, suffer from a scarcity of open-source datasets, hindering the application of LLMs in these specialized areas. Future research can focus on creating domain-specific datasets and fine-tuning LLMs to cater to the unique challenges and intricacies of these fields (Biswas et al., 2020; Sun et al., 2023). Collaboration with domain experts and practitioners is vital to curate relevant data, and fine-tuning LLMs on this data can enhance their effectiveness and ensure better alignment with the specific requirements of each domain, paving the way for LLMs to address real-world challenges (Bubeck et al., 2023) in diverse software engineering domains (Li et al., 2023f).
@@ 8.3. Roadmap
LLMs in software security. The growing impact of LLM4SE offers both unparalleled opportunities and challenges in the domain of software security. On the one hand, LLMs offer promising solutions for automated security audits, compliance verifications, and vulnerability detection. These models can potentially be leveraged for automated code reviews to ensure compliance with industry standards and legal regulations, while also identifying potential security vulnerabilities (Ferrag et al., 2023; Ahmad et al., 2023; Feng and Chen, 2023; Pearce et al., 2023; Deng et al., 2023b; Happe and Cito, 2023). For instance, Ferrag et al. (Ferrag et al., 2023) showcased the efficacy of LLMs in cyber reasoning tasks related to software security. On the other hand, the usage of LLMs introduces novel security concerns. Their complexity makes them susceptible to attacks, demanding novel strategies to fortify the models themselves (Wu et al., 2023c; Rao et al., 2023b; Elizondo, 2023; Ramly, 2023; Deng et al., 2023a; Liu et al., 2023b). As an example, Wu et al. (Wu et al., 2023c) delve into methods to secure LLMs against jailbreak attacks. An intriguing direction for future research lies in enabling LLMs to automatically identify and rectify their own vulnerabilities. Specifically, the focus could be on equipping LLMs to generate self-applied patches to their underlying code, thereby enhancing their inherent security, as opposed to merely implementing application-layer restrictions. Given this landscape, future research should adopt a balanced approach, aiming to exploit LLMs for automating and enhancing existing software security protocols while concurrently developing techniques to secure the LLMs themselves. This dual focus is crucial for fully realizing the potential of LLMs in enhancing the security and compliance assurance of software systems.

### v6

@@ 8.2.3. Enhancing LLM’s Performance in Existing SE Tasks
Tackling domain-specific challenges. Many SE domains, including safety-critical systems and specific industries, suffer from a scarcity of open-source datasets, hindering the application of LLMs in these specialized areas. Future research can focus on creating domain-specific datasets and fine-tuning LLMs to cater to the unique challenges and intricacies of these fields (Biswas et al., 2020; Sun et al., 2023e). Collaboration with domain experts and practitioners is vital to curate relevant data, and fine-tuning LLMs on this data can enhance their effectiveness and ensure better alignment with the specific requirements of each domain, paving the way for LLMs to address real-world challenges (Bubeck et al., 2023) in diverse software engineering domains (Li et al., 2023l).
@@ 8.3. Roadmap
LLMs in software security. The growing impact of LLM4SE offers both unparalleled opportunities and challenges in the domain of software security. On the one hand, LLMs offer promising solutions for automated security audits, compliance verifications, and vulnerability detection. These models can potentially be leveraged for automated code reviews to ensure compliance with industry standards and legal regulations, while also identifying potential security vulnerabilities (Ferrag et al., 2023; Ahmad et al., 2023; Feng and Chen, 2023; Pearce et al., 2023; Deng et al., 2023b; Happe and Cito, 2023). For instance, Ferrag et al. (Ferrag et al., 2023) showcased the efficacy of LLMs in cyber reasoning tasks related to software security. On the other hand, the usage of LLMs introduces novel security concerns. Their complexity makes them susceptible to attacks, demanding novel strategies to fortify the models themselves (Wu et al., 2023e; Rao et al., 2023b; Elizondo, 2023; Ramly, 2023; Deng et al., 2023a; Liu et al., 2023d). As an example, Wu et al. (Wu et al., 2023e) delve into methods to secure LLMs against jailbreak attacks. An intriguing direction for future research lies in enabling LLMs to automatically identify and rectify their own vulnerabilities. Specifically, the focus could be on equipping LLMs to generate self-applied patches to their underlying code, thereby enhancing their inherent security, as opposed to merely implementing application-layer restrictions. Given this landscape, future research should adopt a balanced approach, aiming to exploit LLMs for automating and enhancing existing software security protocols while concurrently developing techniques to secure the LLMs themselves. This dual focus is crucial for fully realizing the potential of LLMs in enhancing the security and compliance assurance of software systems.